OpenStudy (helpmeplease!!!):

From a cruising depth of 75 m below sea lever (BSL) the Nautilus descended to 105 m BSL and then rose to 80 m BSL. What is the total vertical distance the Nuatilus traveled? Another Question - A red kangaroo jumps along 40 km/h. At this rate how long would it take to jump 2 km in minutes?

OpenStudy (helpmeplease!!!):

I have a couple more maths questions: If r = 4, find the value of 3r - 9 4/7 x ? = 16

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

From 1 to 2, the vertical distance is 105 m - 75 m. From 2 to 3, the vertical distance is 105 m - 80 m. Do the two subtractions and add the results.

OpenStudy (mathstudent55):

speed = distance divided by time s = d/t In your case, you know s, the speed of 40 km/h and the distance, 2 km, and you are looking for the time. Solve s = d/t for t: s = d/t Multiply both sides by t st = d Divide both sides by s t = d/s Now substitute the known values for d and s: t = 2 km / (40 km/h) = 0.05 h The result is in hours. To get minutes, multiply hours by 60: t = 0.05 * 60 minutes = 3 minutes

OpenStudy (helpmeplease!!!):

@mathstudent55 thank you soooooo much! that really helped me